# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 5.2 Derivation of Céa estimates
(5.8) $\begin{array}[]{ll}\displaystyle\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{u}}\|_{0,\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|{\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}\\[4.30554pt] \qquad\displaystyle\leq 2(C_{3}^{*}+\widetilde{C}^{*}_{5})(\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{u}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}).\end{array}$
@@
$\begin{array}[]{ll}\displaystyle\|\texttt{e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{\boldsymbol{u}}\|_{0,\Omega}+\|\texttt{e}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}\leq\|{\boldsymbol{\chi}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|{\boldsymbol{\chi}}_{\boldsymbol{u}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{u}}\|_{0,\Omega}\\[4.30554pt] \displaystyle\qquad+\|{\boldsymbol{\chi}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|{\boldsymbol{\chi}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}+\|\xi_{p}\|_{0,\Omega}\\[4.30554pt] \qquad\displaystyle\leq(2C_{3}^{*}+2\widetilde{C}^{*}_{5}+1)(\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{u}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\mbox{\boldmath{$\xi$}}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}),\end{array}$
which combined with the fact that $(\widehat{\boldsymbol{\tau}}_{h},(\widehat{\boldsymbol{v}}_{h},\widehat{{\boldsymbol{\eta}}}_{h}))\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times(\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h})$ and $(\widehat{{\boldsymbol{\psi}}}_{h},\widehat{q}_{h})\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ are arbitrary (see (5.1)), concludes the proof.
@@ 5.3 Rates of convergence
For each $10<m\leq k+$ 1 and for each $\boldsymbol{\tau}\in\mathbb{H}^{m}(\Omega)\cap\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)$ with $\mathbf{div}\,\boldsymbol{\tau}\in\mathbf{H}^{m}(\Omega)$, $\boldsymbol{v}\in\mathbf{H}^{m}(\Omega)$, ${\boldsymbol{\eta}}\in\mathbb{H}^{m}(\Omega)\cap\mathbb{L}^{2}_{\mathrm{skew}}(\Omega)$, ${\boldsymbol{\psi}}\in\mathbf{H}^{m}(\Omega)\cap\mathbf{H}_{\mathrm{N}}(\mathbf{div};\Omega)$ with $\mathop{\rm div}\nolimits\,\boldsymbol{v}\in\mathrm{H}^{m}(\Omega)$, and $q\in\mathrm{H}^{m}(\Omega)$, there holds
(5.9a) $\mathrm{dist}\,\big{(}\boldsymbol{\tau},\mathbb{H}^{\boldsymbol{\sigma}}_{h}\big{)}$ $:=\inf_{\boldsymbol{\tau}_{h}\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}}\|\boldsymbol{\tau}-\boldsymbol{\tau}_{h}\|_{\mathbf{div};\Omega}\lesssim h^{m}\,\Big{\{}\|\boldsymbol{\tau}\|_{m,\Omega}+\|\mathbf{div}\,\boldsymbol{\tau}\|_{m,\Omega}\Big{\}},$
(5.9b) $\mathrm{dist}\,\big{(}\boldsymbol{v},\mathbf{H}^{\boldsymbol{u}}_{h}\big{)}$ $:=\inf_{\boldsymbol{v}_{h}\in\mathbf{H}^{\boldsymbol{u}}_{h}}\|\boldsymbol{v}-\boldsymbol{v}_{h}\|_{0,\Omega}\lesssim h^{m}\,\|\boldsymbol{v}\|_{m,\Omega},$
(5.9c) $\mathrm{dist}\,\big{(}{\boldsymbol{\eta}},\mathbb{H}^{\boldsymbol{\rho}}_{h}\big{)}$ $:=\inf_{{\boldsymbol{\eta}}_{h}\in\mathbb{H}^{\boldsymbol{\rho}}_{h}}\|{\boldsymbol{\eta}}-{\boldsymbol{\eta}}_{h}\|_{0,\Omega}\lesssim h^{m}\,\|{\boldsymbol{\eta}}\|_{m,\Omega},$
(5.9d) $\mathrm{dist}\,\big{(}{\boldsymbol{\psi}},\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\big{)}$ $:=\inf_{{\boldsymbol{\psi}}_{h}\in\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}}\|{\boldsymbol{\psi}}-{\boldsymbol{\psi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}\lesssim h^{m}\,\Big{\{}\|{\boldsymbol{\psi}}\|_{m,\Omega}+\|\mathop{\rm div}\nolimits\,{\boldsymbol{\psi}}\|_{m,\Omega}\Big{\}},$
(5.9e) $\mathrm{dist}\,\big{(}q,\mathrm{H}^{p}_{h}\big{)}$ $:=\inf_{q_{h}\in\mathrm{H}^{p}_{h}}\|q-q_{h}\|_{0,\Omega}\lesssim h^{m}\,\|q\|_{m,\Omega}.$
For (5.9a), (5.9b) and (5.9c) we refer to [33, Th. 2.4], whereas (5.9d) and (5.9e) are provided in [30, Th. 3.6] and [29, Proposition 1.134], respectively. With these steps we are now in position to state the rates of convergence associated with the Galerkin scheme (4.6).
@@
Assume that the hypotheses of Theorem 5.5 hold and let $(\boldsymbol{\sigma},\boldsymbol{u},\boldsymbol{\rho},{\boldsymbol{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ and $(\boldsymbol{\sigma}_{h},\boldsymbol{u}_{h},\boldsymbol{\rho}_{h},{\boldsymbol{\varphi}}_{h},p_{h})\in\mathbb{H}^{\boldsymbol{\sigma}}_{h}\times\mathbf{H}^{\boldsymbol{u}}_{h}\times\mathbb{H}^{\boldsymbol{\rho}}_{h}\times\mathbf{H}^{{\boldsymbol{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ be the unique solutions of the continuous and discrete problems (2.2) and (4.6), respectively. Assume further that $\boldsymbol{\sigma}\in\mathbb{H}^{m}(\Omega)$, $\mathbf{div}\,\boldsymbol{\sigma}\in\mathbf{H}^{m}(\Omega)$, $\boldsymbol{u}\in\mathbf{H}^{m}(\Omega)$, $\boldsymbol{\rho}\in\mathbb{H}^{m}(\Omega)$, ${\boldsymbol{\varphi}}\in\mathbf{H}^{m}(\Omega)$, $\mathop{\rm div}\nolimits\,{\boldsymbol{\varphi}}\in\mathrm{H}^{m}(\Omega)$ and $p\in\mathrm{H}^{m}(\Omega)$, for $11\leq m\leq k+$ 1. Then there exists $C_{\mathrm{rate}}>0$, independent of $h$, such that
$\begin{array}[]{ll}\displaystyle\|\texttt{e}_{\boldsymbol{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{\boldsymbol{u}}\|_{0,\Omega}+\|\texttt{e}_{\boldsymbol{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\boldsymbol{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}\\[4.30554pt] \displaystyle\qquad\leq\,C_{\mathrm{rate}}\,h^{m}\Big{\{}\|\boldsymbol{\sigma}\|_{m,\Omega}\,+\,\|\mathbf{div}\,\boldsymbol{\sigma}\|_{m,\Omega}\,+\,\|\boldsymbol{u}\|_{m,\Omega}\,+\,\|\boldsymbol{\rho}\|_{m,\Omega}\,+\,\|{\boldsymbol{\varphi}}\|_{m,\Omega}\,+\,\|\mathop{\rm div}\nolimits\,{\boldsymbol{\varphi}}\|_{m,\Omega}\,+\,\|p\|_{m,\Omega}\Big{\}}.\end{array}$
Proof 5.8.
The result follows from Céa estimate (5.7) and the approximation properties (5.9).
@@
Similarly to [38], the analysis developed in Sections 2-5 can be adapted to a formulation without the variable $\boldsymbol{\rho}$ ($\boldsymbol{\rho}_{h}$ in the discrete problem), imposing symmetry of $\boldsymbol{\sigma}$ by taking $\boldsymbol{\sigma}\in\mathbb{H}_{\mathrm{sym}}(\mathbf{div};\Omega)\,:=\,\big{\{}\boldsymbol{\tau}\in\mathbb{L}_{\mathrm{sym}}^{2}(\Omega):\ \mathbf{div}\,\boldsymbol{\tau}\in\mathbf{L}^{2}(\Omega)\big{\}}$ and $\mathbb{L}_{\mathrm{sym}}^{2}(\Omega):=\{\boldsymbol{\tau}\in\mathbb{L}^{2}(\Omega):\,\boldsymbol{\tau}=\boldsymbol{\tau}^{\tt t}\}$, utilizing results from [38, Section 2.2] ([38, Section 4.1] for the discrete problem), and adapting the strategy used in, e.g., [38, Sections 3 and 4].
@@ 6 A posteriori error estimates
$\mathcal{E}_{h}(\Omega)\,:=\,\big{\{}\,E\in\mathcal{E}_{h}\,:\;E\subseteq\Omega\,\big{\}},\quad\mathcal{E}_{h}(\Gamma_{\mathrm{D}})\,:=\,\big{\{}\,E\in\mathcal{E}_{h}\,:\;E\subseteq\Gamma_{\mathrm{D}}\,\big{\}},{\quad\hbox{and}\quad}\mathcal{E}_{h}(\Gamma_{\mathrm{N}})\,:=\,\big{\{}\,E\in\mathcal{E}_{h}\,:\;E\subseteq\Gamma_{\mathrm{N}}\,\big{\}}\,.$
@@
$\mathrm{curl}\,(\psi):=\Big{(}\frac{\partial\psi}{\partial x_{2}}\,,\,-\frac{\partial\psi}{\partial x_{1}}\Big{)}^{\tt t}\,,\quad\mathrm{rot}\,(\boldsymbol{v}):=\frac{\partial v_{2}}{\partial x_{1}}-\frac{\partial v_{1}}{\partial x_{2}}\,,\quad\mathbf{curl}\,(\boldsymbol{v})=\begin{pmatrix}\mathrm{curl}\,(v_{1})^{\mathrm{t}}\\[4.30554pt] \mathrm{curl}\,(v_{2})^{\mathrm{t}}\end{pmatrix}{\quad\hbox{and}\quad}\underline{\mathbf{curl}\,}(\boldsymbol{\tau})=\begin{pmatrix}\mathrm{rot}\,(\boldsymbol{\tau}_{1})\\ \mathrm{rot}\,(\boldsymbol{\tau}_{2})\end{pmatrix}.$
@@
In what follows, we will assume that the hypotheses of Theorems 3.12 and 4.1 are satisfied. Let $\boldsymbol{\sigma}_{h}$, $\boldsymbol{u}_{h}$, $\boldsymbol{\rho}_{h}$, ${\boldsymbol{\varphi}}_{h}$, $p_{h}$ denote the FE solutions of (4.6). We define the residual-based and fully computable local contributions to the error estimator $\Xi_{K}^{2}$, defined as the sum of $\Xi_{s,K}^{2}$ and $\Xi_{f,K}^{2}$, where $\Xi_{s,K}$ and $\Xi_{f,K}$ pertain to the solid (mixed elasticity) and fluid (mixed Darcy) components, respectively:
$\Xi_{s,K}^{2}$ $:=\|\boldsymbol{f}+\mathbf{div}\boldsymbol{\sigma}_{h}\|_{0,K}^{2}+\|\boldsymbol{\sigma}_{h}-\boldsymbol{\sigma}_{h}^{\tt t}\|^{2}_{0,K}+h_{K}^{2}\|\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}-\boldsymbol{\nabla}\boldsymbol{u}_{h}+\boldsymbol{\rho}_{h}\|_{0,K}^{2}$
$\quad+h_{K}^{2}\|\underline{\mathbf{curl}}(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\|_{0,K}^{2}+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Omega)}h_{E}\|[\![(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}]\!]\|_{0,E}^{2}$
(6.1a) $\quad+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|(\mathcal{C}^{-1}\boldsymbol{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbb{I}+\boldsymbol{\rho}_{h})\boldsymbol{s}-\frac{\mathrm{d}\boldsymbol{u}_{\mathrm{D}}}{\mathrm{d}\boldsymbol{s}}\|_{0,E}^{2}+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|\boldsymbol{u}_{\mathrm{D}}-\boldsymbol{u}_{h}\|_{0,E}^{2},$
$\Xi_{f,K}^{2}$ $:=\|-g+c_{0}p_{h}+\dfrac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\boldsymbol{\sigma}_{h}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}p_{h}-\mathop{\rm div}\nolimits{\boldsymbol{\varphi}}_{h}\|_{0,K}^{2}+h_{K}^{2}\|\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}{\boldsymbol{\varphi}}_{h}-\nabla p_{h}\|_{0,K}^{2}$
$\quad+h_{K}^{2}\|\mathrm{rot}\,(\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}{\boldsymbol{\varphi}}_{h})\|_{0,K}^{2}+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Omega)}h_{E}\|[\![(\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}{\boldsymbol{\varphi}}_{h})\cdot\boldsymbol{s}]\!]\|_{0,E}^{2}$
(6.1b) $\quad+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|\kappa(\boldsymbol{\sigma}_{h},p_{h})^{-1}{\boldsymbol{\varphi}}_{h}\cdot\boldsymbol{s}-\frac{\mathrm{d}p_{\mathrm{D}}}{\mathrm{d}\boldsymbol{s}}\|_{0,E}^{2}+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|p_{\mathrm{D}}-p_{h}\|_{0,E}^{2}.$

### v2

@@ 5.2 Derivation of Céa estimates
$\begin{array}[]{ll}\displaystyle\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|{\bm{\chi}}_{{\bm{u}}}\|_{0,\Omega}+\|{\bm{\chi}}_{\bm{\rho}}\|_{0,\Omega}+\|{\bm{\chi}}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}\\[4.30554pt] \qquad\displaystyle\leq 2(C_{3}^{*}+\widetilde{C}^{*}_{5})(\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\bm{\xi}_{{\bm{u}}}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\rho}}\|_{0,\Omega}+\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}).\end{array}$ (5.8)
@@
$\|\texttt{e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{{\bm{u}}}\|_{0,\Omega}+\|\texttt{e}_{\bm{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}$
$\quad\leq\|{\bm{\chi}}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|{\bm{\chi}}_{{\bm{u}}}\|_{0,\Omega}+\|\bm{\xi}_{{\bm{u}}}\|_{0,\Omega}$
$\qquad+\|{\bm{\chi}}_{\bm{\rho}}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\rho}}\|_{0,\Omega}+\|{\bm{\chi}}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\chi_{p}\|_{0,\Omega}+\|\xi_{p}\|_{0,\Omega}$
$\quad\leq(2C_{3}^{*}+2\widetilde{C}^{*}_{5}+1)(\|\bm{\xi}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\bm{\xi}_{{\bm{u}}}\|_{0,\Omega}+\|\bm{\xi}_{\bm{\rho}}\|_{0,\Omega}+\|\bm{\xi}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\xi_{p}\|_{0,\Omega}),$
which combined with the fact that $(\widehat{\bm{\tau}}_{h},(\widehat{\mathbf{v}}_{h},\widehat{{\bm{\eta}}}_{h}))\in\mathbb{H}^{\bm{\sigma}}_{h}\times(\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h})$ and $(\widehat{{\bm{\psi}}}_{h},\widehat{q}_{h})\in\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ are arbitrary (see (5.1)), concludes the proof. ∎
@@ 5.3 Rates of convergence
For each $10<m\leq k+$ 1 and for each $\bm{\tau}\in\mathbb{H}^{m}(\Omega)\cap\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)$ with $\mathbf{div}\,\bm{\tau}\in\mathbf{H}^{m}(\Omega)$, $\mathbf{v}\in\mathbf{H}^{m}(\Omega)$, ${\bm{\eta}}\in\mathbb{H}^{m}(\Omega)\cap\mathbb{L}^{2}_{\mathrm{skew}}(\Omega)$, ${\bm{\psi}}\in\mathbf{H}^{m}(\Omega)\cap\mathbf{H}_{\mathrm{N}}({\mathop{\rm div}\nolimits};\Omega)$ with $\mathop{\rm div}\nolimits\,\mathbf{v}\in\mathrm{H}^{m}(\Omega)$, and $q\in\mathrm{H}^{m}(\Omega)$, there holds
$\mathrm{dist}\,\big{(}\bm{\tau},\mathbb{H}^{\bm{\sigma}}_{h}\big{)}$ $:=\inf_{\bm{\tau}_{h}\in\mathbb{H}^{\bm{\sigma}}_{h}}\|\bm{\tau}-\bm{\tau}_{h}\|_{\mathbf{div};\Omega}\lesssim h^{m}\,\Big{\{}\|\bm{\tau}\|_{m,\Omega}+\|\mathbf{div}\,\bm{\tau}\|_{m,\Omega}\Big{\}},$ (5.9a)
$\mathrm{dist}\,\big{(}\mathbf{v},\mathbf{H}^{{\bm{u}}}_{h}\big{)}$ $:=\inf_{\mathbf{v}_{h}\in\mathbf{H}^{{\bm{u}}}_{h}}\|\mathbf{v}-\mathbf{v}_{h}\|_{0,\Omega}\lesssim h^{m}\,\|\mathbf{v}\|_{m,\Omega},$ (5.9b)
$\mathrm{dist}\,\big{(}{\bm{\eta}},\mathbb{H}^{\bm{\rho}}_{h}\big{)}$ $:=\inf_{{\bm{\eta}}_{h}\in\mathbb{H}^{\bm{\rho}}_{h}}\|{\bm{\eta}}-{\bm{\eta}}_{h}\|_{0,\Omega}\lesssim h^{m}\,\|{\bm{\eta}}\|_{m,\Omega},$ (5.9c)
$\mathrm{dist}\,\big{(}{\bm{\psi}},\mathbf{H}^{{\bm{\varphi}}}_{h}\big{)}$ $:=\inf_{{\bm{\psi}}_{h}\in\mathbf{H}^{{\bm{\varphi}}}_{h}}\|{\bm{\psi}}-{\bm{\psi}}_{h}\|_{\mathop{\rm div}\nolimits;\Omega}\lesssim h^{m}\,\Big{\{}\|{\bm{\psi}}\|_{m,\Omega}+\|\mathop{\rm div}\nolimits\,{\bm{\psi}}\|_{m,\Omega}\Big{\}},$ (5.9d)
$\mathrm{dist}\,\big{(}q,\mathrm{H}^{p}_{h}\big{)}$ $:=\inf_{q_{h}\in\mathrm{H}^{p}_{h}}\|q-q_{h}\|_{0,\Omega}\lesssim h^{m}\,\|q\|_{m,\Omega}.$ (5.9e)
For (5.9a), (5.9b) and (5.9c) we refer to [34, Th. 2.4], whereas (5.9d) and (5.9e) are provided in [31, Th. 3.6] and [29, Proposition 1.134], respectively. With these steps we are now in a position to state the rates of convergence associated with the Galerkin scheme (4.5).
@@
Assume that the hypotheses of Theorem 5.3 hold and let $(\bm{\sigma},{\bm{u}},\bm{\rho},{\bm{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$ and $(\bm{\sigma}_{h},{\bm{u}}_{h},\bm{\rho}_{h},{\bm{\varphi}}_{h},p_{h})\in\mathbb{H}^{\bm{\sigma}}_{h}\times\mathbf{H}^{{\bm{u}}}_{h}\times\mathbb{H}^{\bm{\rho}}_{h}\times\mathbf{H}^{{\bm{\varphi}}}_{h}\times\mathrm{H}^{p}_{h}$ be the unique solutions of the continuous and discrete problems (2.2) and (4.5), respectively. Assume further that $\bm{\sigma}\in\mathbb{H}^{m}(\Omega)$, $\mathbf{div}\,\bm{\sigma}\in\mathbf{H}^{m}(\Omega)$, ${\bm{u}}\in\mathbf{H}^{m}(\Omega)$, $\bm{\rho}\in\mathbb{H}^{m}(\Omega)$, ${\bm{\varphi}}\in\mathbf{H}^{m}(\Omega)$, $\mathop{\rm div}\nolimits\,{\bm{\varphi}}\in\mathrm{H}^{m}(\Omega)$ and $p\in\mathrm{H}^{m}(\Omega)$, for $11\leq m\leq k+$ 1. Then there exists $C_{\mathrm{rate}}>0$, independent of $h$, such that
$\begin{array}[]{ll}\displaystyle\|\texttt{e}_{\bm{\sigma}}\|_{\mathbf{div};\Omega}+\|\texttt{e}_{{\bm{u}}}\|_{0,\Omega}+\|\texttt{e}_{\bm{\rho}}\|_{0,\Omega}+\|\texttt{e}_{{\bm{\varphi}}}\|_{\mathop{\rm div}\nolimits;\Omega}+\|\texttt{e}_{p}\|_{0,\Omega}\\[4.30554pt] \displaystyle\quad\leq\ C_{\mathrm{rate}}\,h^{m}\Big{\{}\|\bm{\sigma}\|_{m,\Omega}+\|\mathbf{div}\,\bm{\sigma}\|_{m,\Omega}+\|{\bm{u}}\|_{m,\Omega}+\|\bm{\rho}\|_{m,\Omega}+\|{\bm{\varphi}}\|_{m,\Omega}+\|\mathop{\rm div}\nolimits\,{\bm{\varphi}}\|_{m,\Omega}+\|p\|_{m,\Omega}\Big{\}}.\end{array}$
Proof.
The result follows from Céa estimate (5.7) and the approximation properties (5.9). ∎
@@
Similarly to [39], the analysis developed in Sections 2-5 can be adapted to a formulation without the variable $\bm{\rho}$ ($\bm{\rho}_{h}$ in the discrete problem), imposing symmetry of $\bm{\sigma}$ by taking $\bm{\sigma}\in\mathbb{H}_{\mathrm{sym}}(\mathbf{div};\Omega)\,:=\,\big{\{}\bm{\tau}\in\mathbb{L}_{\mathrm{sym}}^{2}(\Omega):\ \mathbf{div}\,\bm{\tau}\in\mathbf{L}^{2}(\Omega)\big{\}}$ and $\mathbb{L}_{\mathrm{sym}}^{2}(\Omega):=\{\bm{\tau}\in\mathbb{L}^{2}(\Omega):\,\bm{\tau}=\bm{\tau}^{\tt t}\}$, utilizing results from [39, Section 2.2] ([39, Section 4.1] for the discrete problem), and adapting the strategy used in, e.g., [39, Sections 3 and 4].
@@ 6 A posteriori error estimates
${\color[rgb]{0.22,0.45,0.70}\definecolor[named]{pgfstrokecolor}{rgb}{0.22,0.45,0.70}\mathcal{E}_{h}(\star)\,:=\,\big{\{}\,E\in\mathcal{E}_{h}\,:\;E\subseteq\star\,\big{\}},\quad\star\in\{\Omega,\Gamma_{\mathrm{D}},\Gamma_{\mathrm{N}}\}}.$
@@
$\mathrm{curl}\,(\psi):=\Big{(}\frac{\partial\psi}{\partial x_{2}}\,,\,-\frac{\partial\psi}{\partial x_{1}}\Big{)}^{\tt t}\,,\quad\mathrm{rot}\,(\mathbf{v}):=\frac{\partial v_{2}}{\partial x_{1}}-\frac{\partial v_{1}}{\partial x_{2}}\,,\quad\mathbf{curl}\,(\mathbf{v})=\begin{pmatrix}\mathrm{curl}\,(v_{1})^{\mathrm{t}}\\[4.30554pt] \mathrm{curl}\,(v_{2})^{\mathrm{t}}\end{pmatrix},$
${\quad\hbox{and}\quad}\underline{\mathbf{curl}\,}(\bm{\tau})=\begin{pmatrix}\mathrm{rot}\,(\bm{\tau}_{1})\\ \mathrm{rot}\,(\bm{\tau}_{2})\end{pmatrix}.$
@@
In what follows, we will assume that the hypotheses of Theorems 3.8 and 4.1 are satisfied. Let $\bm{\sigma}_{h}$, ${\bm{u}}_{h}$, $\bm{\rho}_{h}$, ${\bm{\varphi}}_{h}$, $p_{h}$ denote the FE solutions of (4.5). We define the residual-based and fully computable local contributions to the error estimator $\Xi_{K}^{2}$, defined as the sum of $\Xi_{s,K}^{2}$ and $\Xi_{f,K}^{2}$, where $\Xi_{s,K}$ and $\Xi_{f,K}$ pertain to the solid (mixed elasticity) and fluid (mixed Darcy) components, respectively:
$\Xi_{s,K}^{2}$ $:=\|\bm{f}+\mathbf{div}\bm{\sigma}_{h}\|_{0,K}^{2}+\|\bm{\sigma}_{h}-\bm{\sigma}_{h}^{\tt t}\|^{2}_{0,K}+h_{K}^{2}\|\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}-\bm{\nabla}{\bm{u}}_{h}+\bm{\rho}_{h}\|_{0,K}^{2}$
$\quad+h_{K}^{2}\|\underline{\mathbf{curl}}(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\|_{0,K}^{2}+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|{\bm{u}}_{\mathrm{D}}-{\bm{u}}_{h}\|_{0,E}^{2}$
$\quad+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Omega)}h_{E}\|[\![(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}]\!]\|_{0,E}^{2}$
$\quad+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|(\mathcal{C}^{-1}\bm{\sigma}_{h}+\dfrac{\alpha}{d\lambda+2\mu}p_{h}\mathbf{I}+\bm{\rho}_{h})\bm{s}-\frac{\mathrm{d}{\bm{u}}_{\mathrm{D}}}{\mathrm{d}\bm{s}}\|_{0,E}^{2},$ (6.1a)
$\Xi_{f,K}^{2}$ $:=\|c_{0}p_{h}+\dfrac{\alpha}{d\lambda+2\mu}\mathop{\rm tr}\bm{\sigma}_{h}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}p_{h}-\mathop{\rm div}\nolimits{\bm{\varphi}}_{h}-g\|_{0,K}^{2}+h_{K}^{2}\|{\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}{\bm{\varphi}}_{h}-\nabla p_{h}\|_{0,K}^{2}$
$\quad+h_{K}^{2}\|\mathrm{rot}\,({\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}{\bm{\varphi}}_{h})\|_{0,K}^{2}+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Omega)}h_{E}\|[\![({\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}{\bm{\varphi}}_{h})\cdot\bm{s}]\!]\|_{0,E}^{2}$
$\quad+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|{\bm{\kappa}}(\bm{\sigma}_{h},p_{h})^{-1}{\bm{\varphi}}_{h}\cdot\bm{s}-\frac{\mathrm{d}p_{\mathrm{D}}}{\mathrm{d}\bm{s}}\|_{0,E}^{2}+\sum_{E\in\partial K\cap\mathcal{E}_{h}(\Gamma_{\mathrm{D}})}h_{E}\|p_{\mathrm{D}}-p_{h}\|_{0,E}^{2}.$ (6.1b)